OpenStudy (anonymous):

Geometry question concerning distance from point A to point B. Question is the first reply. Find the distance between points M(6,16) and Z(-1,14) to the nearest tenth

OpenStudy (anonymous):

Find the distance between points M(6,16) and Z(-1,14) to the nearest tenth

OpenStudy (anonymous):

think method first from 6 to -1 is 7 units from 16 to 14 is 2 units pythagoras gives \[\sqrt{7^2+2^2}\]

OpenStudy (anonymous):

i found it easier with the thinking method after i understood it. does a thinking method apply here as well?

OpenStudy (anonymous):

ah thats what youre doing

OpenStudy (anonymous):

im sorry your comment just showed up

OpenStudy (deepolisnoob):

For this problem, you will need to use the Pythagorean theorem. On the x-axis, the points move 7 units, while on the y-axis, the points move 2 units. Knowing that one side of your triangle is 7 units long and the other side is 2 units long, you can square 7 and 2, then add them, being equal 49+4=53. Then, you need to find the square root of 53. The equation is a^2+b^2=c^2, c^2 being the side you want to find.

OpenStudy (anonymous):

i know Pythagorean theorem. that will make this a lot easier.

OpenStudy (anonymous):

yes you have a right triangle the base has length 7 (from 6 to -1) the height has length 2 (from 16 to 14) the hypotenuse is the distance you are looking for via pythagoras it is \[d^2=a^2+b^2\\ d^2=7^2+2^2\\ d=\sqrt{7^2+2^2}\]

OpenStudy (anonymous):

let me try it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

7.3 units

OpenStudy (anonymous):

rounded to the tenths place

OpenStudy (anonymous):

let me check it looks right did you get \[\sqrt{53}\] first?

OpenStudy (anonymous):

yeah 7.3 rounded

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then took the square root of 53

OpenStudy (anonymous):

got 7.28

OpenStudy (anonymous):

rounded

OpenStudy (anonymous):

you are right